# Bayesian Low-Rank LeArning (Bella): A Practical Approach to Bayesian Neural Networks

### v1

@@ 5.3 Performance on Out-Of-Disrtibution (OOD) Datasets
Table 4: Out-of-distribution performance of our method Bella on the CIFAR-10-C benchmark. (Recall Bayesian approximation are from $n=5$ particles for Ensemble and SVGD).
5.4 Comparing Uncertainty Estimations
Bayesian models capable of providing a theoretical basis for measuring model uncertainty. Also known as epistemic uncertainty, refers to uncertainty stemming from limitations in our knowledge or understanding of the underlying data generating process or the model itself. One of the ways to quantify model uncertainty is through mutual information estimates, following  [47].
Figure 2: Evaluation of uncertainty estimations using Mutual Information on CAMELYON17 and CIFAR-10-C datasets. $\uparrow$ MI for Misclassified Examples is better—denoted by the distribution shifting $\rightarrow$. In contrast, $\downarrow$ MI for Correctly Classified Examples is better—denoted by the distribution shifting $\leftarrow$ (Notably, Single base models are not able to capture epistemic uncertainty [47].)
5.5 Robustness against Adversarial Examples
In this section, we examine the resilience of our proposed Bella against adversarial attacks, specifically employing the $L_{\infty}$ Fast Gradient Sign Method (FGSM) across various attack budgets as detailed in Figure 3. This analysis aims to benchmark the robustness of our method in comparison to traditional models under adversarial conditions. We employ the robustness benchmark [48] to deploy the attack on CIFAR-10 test set and report results in Figure 3.
Figure 3: Comparison of model robustness to $L_{\infty}$ FGSM adversarial attacks across varied attack budgets. The evaluation is on the CIFAR-10 dataset. Bella achieves comparable performance to current, baseline SVGD and Ensemble implementations whilst demonstrating significantly better robustness (higher accuracy) than a Single baseline model.
5.6 Ablation Studies
This section undertakes a series of ablation studies to examine the effects of various components within Bella. Our analysis includes an exploration of the training costs associated with different ranks and their consequent influence on model performance. Given that Bella incorporates multiple parameter particles, we also delve into how varying the number of these particles affects Bella’s efficacy. Additionally, we explore the application of low-rank adapters across different layers and assess their impact. Further details on other studies are in Appendix D. We show in Table 6 that we achieve state-of-the-art performance on CAMELYON17.
Ablations on rank $r$As outlined in Section 4, we substitute the network’s extensive full matrix with low-rank matrices, which are defined by the ‘rank’ parameter ($r$). This section aims to assess how this parameter influences Bella’s performance.
To comprehensively demonstrate the rank’s impact, we perform our analysis on the demanding large-scale CAMELYON17 dataset. The findings, illustrated in Figure 4(a), shed light on how the number of ranks affects performance. Utilizing a small rank significantly reduces the parameter space, but this constriction limits Bella’s ability to learn effectively and achieve optimal performance. Conversely, enhancing the rank to 4 markedly boosts Bella’s efficiency. It’s observed that performance plateaus at a rank of 16, indicating a saturation point.
Figure 4: The impact of ranks on CAMELYON17 performance (a), as well as the impact of the number of parameter particles on CIFAR-10 (b) on Bella performance.
Ablations on the number of particles $n$In this section, we delve into the influence of the quantity of parameter particles on the performance of Bella. The findings, depicted in Figure 4(b), reveal an improvement in Bella’s performance with an increase in the number of parameter particles.
Figure 5: Evaluation on CAMELYON17 and ImageNet vs. Full SVGD: Bella achieves comparable performance with significantly fewer parameters.
The starting point of the plot in Figure 5 associated with the CAMELYON17 dataset is for Bella with $r=2$ (consisting of $5$ particles), and the end point is associated with full SVGD (SVGD base) with the same number of particles. As $r$ increases, the accuracy of Bellas improves and after a while becomes plateau around $95.08$, comparable to the accuracy of full SVGD ($95.21$). This underscores the efficiency and advantages of our proposed Bella. Despite utilizing a substantially smaller pool of trainable parameters (around 0.3% for $r=4$) compared to full SVGD models, Bella demonstrates performance on par with these more parameter-intensive alternatives.
5.7 Generalization to a Visual Question Answer (VQA) Task
In this section, we extend the application of our Bella to another challenging vision task, Visual Question Answering (VQA), as detailed in [49]. We leverage the state-of-the-art, pre-trained, large multi-modal model LlaVA  [13, 50] for this purpose. Utilizing LlaVA transforms VQA into a process where an image and a natural-language question are inputs, and the model generates a free-form, open-ended text answer. Answering questions in VQA requires various intelligent capabilities, including not only image recognition but also complex reasoning. For this task, we employ VQA v2 [49] dataset containing 204,721 images, more than 1 Billion (1B) questions and 10B ground-truth answers in total. There are three main types of answers: Yes/No, a Number, and Other.
Model. In our experiments, we employed our proposed Bella on top of LlaVA-1.5-7B [13, 50]. Further details about the dataset, model and metrics are deferred to Appendix F.
Results (Performance).  The primary outcomes for Yes/No and Number answer queries are detailed in Table 5, respectively. We chose these question types to ensure a fair comparison and avoid semantic mismatches in open-ended answers in Others. A distinctive feature of our Bella is its reduced uncertainty estimates to accurate predictions, coupled with increased uncertainty estimates to incorrect ones. Moreover, it surpasses the Single base model in terms of Accuracy and Exact Match metrics.
Notably, our method is efficient, particularly in contrast to the resource-intesive baseline Bayesian models tailored for this task (e.g. [51, 52, 6]). The computational requirements for using these baselines render the application of full SVGD or ensemble alternatives impractical, thereby highlighting the impact on practical applications by harnessing the effectiveness of Bayesian models with Bella.
VQA: Evaluation of Yes/No Questions
Models Entropy of Corrects ($\downarrow$) Entropy of Incorrect ($\uparrow$) Accuracy ($\uparrow$) Exact Match ($\uparrow$)
Single base 0.3336 0.5920 91.59 86.74
     Ensemble Bella 0.3438 0.5935 91.20 86.21
SVGD Bella 0.3245 0.5950 92.46 87.83
VQA: Evaluation of Number Questions
Single base 0.4148 0.9911 58.69 49.26
Ensemble Bella 0.4248 0.9929 57.86 48.68
SVGD Bella 0.4059 0.9816 60.19 50.99
Results (Model Uncertainty and Human Confidence). Further, we investigate the relationship between model uncertainty attained from Bella and human confidence, facilitated by multiple annotators in the VQA dataset. For this, we measure the correlation between the model disagreement by measuring the predictive entropy vs the human annotations of the answers.
(a) Yes/No questions.
(b) Number questions.
In Fig. 6 we shows the correlation between the entropy of the model outputs and human confidence levels. By setting a specific threshold for the model’s entropy, we effectively bifurcate our predictions into two distinct categories: those with lower entropy fall into the ‘Low Entropy’ group, signalling reduced uncertainty within the model’s assessments, while predictions surpassing this entropy threshold are allocated to the ‘High Entropy’ segment, indicative of greater uncertainty. Intriguingly, our observations reveal a consistent negative correlation between the model’s entropy levels and human confidence across different types of queries, such as Yes/No and Number questions. This pattern suggests that the model’s entropy can serve as a reliable gauge of certainty, mirroring human judgment in its response to varying levels of uncertainty.
6 Conclusion
A Practical Approach to Bayesian Neural Networks
Overview of Materials in the Appendices
  1. 1.
    Diversity measures between SVGD Bella and its SVGD base model (Appendix A)
  2. 2.
  3. 3.
    Detailed uncertainty measured on corrupted CIFAR-10-C dataset (Appendix C)
  4. 4.
    Additional study on weight averaging (Appendix D)
  5. 5.
    Impact of the pushing parameter on the robustness of models (Appendix E)
  6. 6.
    Detailed information about VQA task (Appendix F)
  7. 7.
  8. 8.
    Detailed information regarding datasets and hyper-parameters utilized in the paper (Appendix H)
Appendix A Diversity Measures
Appendix B Generalization of CIFAR-10 Models on OOD dataset (STL-10)
In this section, we evaluate the generalization of networks trained on CIFAR-10 on a similar datasets, which is STL-10.
@@ Appendix C Measure Uncertainty on Corrupted CIFAR-10-C Dataset
Appendix D Additional Studies
@@
Recent research has increasingly explored the concept of weight averaging, a technique where parameters from multiple models are combined to create a unified model that may offer superior predictive performance [53, 54, 55]. This approach is grounded in the theory that models starting from the same pre-trained state tend to have a linear loss landscape, making averaging a viable strategy. For a detailed mathematical discussion, see [56]. Corroborating this, Figure 10 features density plots and Table 7 that illustrate the predictive entropy for samples that are correctly and incorrectly classified by the CIFAR-10 and CIFAR-100 datasets using the ‘soup’ models. These results show that the performance of Bella models, post-averaging, is comparable to that of more computationally intensive baseline models. This suggests that Bella models, despite having fewer trainable parameters, are capable of achieving the same level of performance as their more complex counterparts.
@@
Table 6: Choosing different layers for fine-tuning with Bella, with $r=4,n=5$ on CAMELYON17 achieves state-of-the-art performance [11].
Ablations on fine-tuned layers.  In our paper, we propose the concept of substituting the network’s complete weight matrices with low-rank adapters to approximate BNNs. This section is dedicated to assessing the effectiveness of these low-rank adapters across various layer configurations within CLIP, utilizing CAMELYON17. It should be noted that CLIP consists of 12 ResidualAttentionBlocks, each comprising Fully Connected (FC) and Projection (Proj.) linear layers, which we aim to modify for our Bella. The data, presented in Table 6, reveals interesting performance of our Bella as it surpasses the SOTA in one of the experiments. This highlights the adaptability of our proposed approach, which does not necessitate extensive fine-tuning to select specific layers for optimal performance.
Model Name CIFAR-10 CIFAR-100
Accuracy PE Correct PE Incorrect Accuracy PE Correct PE Incorrect
Ensemble Bella soup 95.74 0.0597 0.6063 80.82 0.3461 1.2261
Ensemble base soup 97.42 0.0986 0.6917 86.42 0.3025 1.1148
SVGD Bella soup 96.32 0.0771 0.6988 83.59 0.2935 1.1044
SVGD base soup 97.62 0.0438 0.5528 86.73 0.2609 1.0426
Ensemble base 97.54 0.1125 0.5947 86.89 0.3738 1.0677
SVGD base 97.87 0.0572 0.4692 87.32 0.3256 1.0015
Ensemble Bella 97.58 0.0685 0.5098 86.50 0.4100 1.1218
SVGD Bella 97.95 0.0603 0.5193 88.09 0.4000 1.0915
@@
Comparison of SVGD Bella and VBL.
In this section, we compare the performance of SVGD Bella and VBL in distinguishing between correctly and misclassified samples. Specifically, we assess their ability to assign high entropy to misclassified samples and low entropy to correctly classified ones. Figure 11 shows that SVGD Bella outperforms VBL in this regard.
Appendix E Impact of $\gamma$ values on Robustness.
@@
Appendix F Detailed information regarding VQA task
@@
F.1 Task Definition and Experiment Setup
Task. Visual Question Answering (VQA) [49] is a free-form and open-ended task, taking as input an image and a natural-language question about the image and producing a natural-language answer as the output. Questions in VQA require various intelligence capabilities to answer, including image recognition and object detection, as well as reasoning like commonsense reasoning. Below is an overview of the VQA dataset utilized in our paper:
  • VQA v2 [42]: This dataset contains 204,721 images, more than 1 Billion (1B) questions, and 10B ground truth answers in total. There are three main types of answers: Yes/No, Number, and Other. The evaluation set contains 80,541, 28,134 and 105,679 questions for Yes/No, Number, and Other respectively.
@@ F.1 Task Definition and Experiment Setup
Models. In our experiments, we applied Bella on top of LLaVA [13, 50] to address the VQA task. Specifically, we utilized the LLaVA-1.5 7B model, an improved version of the original LLaVA with superficial modifications. We followed the public training method using a deepspeed codebase.
@@ F.2 Experimental Results
Results. The performance of different single models and ensembles are presented in Table 5, Table 10. We ensembled two Bella models by simply averaging two softmax output scores. Given the model similarity matrix in Table 9, some observations and conclusions can be made from these tables.
@@
Models Correct Entropy Incorrect Entropy Accuracy Exact Match
Single base 0.5782 1.4826 70.00 57.17
Ensemble Bella 0.6079 1.5457 67.05 55.26
SVGD Bella 0.5359 1.4602 68.38 56.68
@@
Figure 13: In comparing LLaVA-VQA predictions on ‘Other’ questions: SVGD Bella outperforms Ensemble Bella, even with more extended predictions. In one case, both Bella err, but with notable human annotator disagreement. Interestingly, SVGD Bella shows greater uncertainty (MI), unlike the confident Ensemble Bella (MI=0). Given that 2 out of 10 annotators chose "Brown," SVGD Bella’s answer is deemed acceptable. In another case, SVGD Bella correctly identifies "Giraffe" albeit singularly, mirroring a minor error also seen among human annotators.
@@
Appendix G RBF kernel computation in SVGD
The trace term can be written as
$\mathrm{Tr}[\mathbf{A}_{i}\mathbf{A}_{i}^{\top}\mathbf{B}_{i}^{\top}\mathbf{B}_{i}]+\mathrm{Tr}[\mathbf{A}_{i}\mathbf{A}_{i}^{\top}\mathbf{B}_{i}^{\top}\mathbf{B}_{i}]-\mathrm{Tr}[\mathbf{A}_{j}\mathbf{A}_{i}^{\top}\mathbf{B}_{i}^{\top}\mathbf{B}_{j}]-\mathrm{Tr}[\mathbf{A}_{i}\mathbf{A}_{j}^{\top}\mathbf{B}_{j}^{\top}\mathbf{B}_{i}].$
Appendix H Detailed Parameters and Datasets

### v2

@@ Appendix B Generalization of CIFAR-10 Models on OOD dataset (STL-10)
Table 5: Out-of-distribution performance (in terms of accuracy) of our method Bella on the CIFAR-10-C benchmark. (Recall Bayesian approximation are from $n=5$ particles for Ensemble and SVGD).
@@ Appendix C Measure Uncertainty on Corrupted CIFAR-10-C Dataset
Model Name CIFAR-10 CIFAR-100
Accuracy $\uparrow$ PE Correct $\downarrow$ PE Incorrect $\uparrow$ Accuracy $\uparrow$ PE Correct $\downarrow$ PE Incorrect $\uparrow$
Ensemble Bella soup 95.74 0.0597 0.6063 80.82 0.3461 1.2261
Ensemble base soup 97.42 0.0986 0.6917 86.42 0.3025 1.1148
SVGD Bella soup 96.32 0.0771 0.6988 83.59 0.2935 1.1044
SVGD base soup 97.62 0.0438 0.5528 86.73 0.2609 1.0426
Ensemble base 97.54 0.1125 0.5947 86.89 0.3738 1.0677
Ensemble Bella 97.58 0.0685 0.5098 86.50 0.4100 1.1218
SVGD base 97.87 0.0572 0.4692 87.32 0.3256 1.0015
SVGD Bella 97.95 0.0603 0.5193 88.09 0.4000 1.0915
Appendix D Generalization Performance of Models on CIFAR-10-C Task
The generalization performance of different models in terms of accuracy across various noise types (with a maximum severity level of 5) is depicted in Table 5. The results show that Bella SVGD performs comparably to its base counterpart.
Appendix E Additional Comparison Studies
@@
Recent research has increasingly explored the concept of weight averaging, a technique where parameters from multiple models are combined to create a unified model that may offer superior predictive performance (Matena and Raffel 2022; Wortsman et al. 2022a, b). This approach is grounded in the theory that models starting from the same pre-trained state tend to have a linear loss landscape, making averaging a viable strategy. For a detailed mathematical discussion, see (Rame et al. 2022). Corroborating this, Figure 10 features density plots and Table 6 that illustrate the predictive entropy for samples that are correctly and incorrectly classified by the CIFAR-10 and CIFAR-100 datasets using the ‘soup’ models. These results show that the performance of Bella models, post-averaging, is comparable to that of more computationally intensive baseline models. This suggests that Bella models, despite having fewer trainable parameters, are capable of achieving the same level of performance as their more complex counterparts.
@@
Table 7: Choosing different layers for fine-tuning with Bella, with $r=4,n=5$ on CAMELYON17 achieves state-of-the-art performance (Challenge 2024) (see highlighted row).
@@
Ablations on fine-tuned layers.  In our paper, we propose the concept of substituting the network’s complete weight matrices with low-rank adapters to approximate BNNs. This section is dedicated to assessing the effectiveness of these low-rank adapters across various layer configurations within CLIP, utilizing CAMELYON17. It should be noted that CLIP consists of 12 ResidualAttentionBlocks, each comprising Fully Connected (FC) and Projection (Proj.) linear layers, which we aim to modify for our Bella. The data, presented in Table 7, reveals an interesting performance outcome from our Bella as it surpasses the SOTA in one of the experiments. This highlights the adaptability of our proposed approach, which does not necessitate extensive fine-tuning to select specific layers for optimal performance.
Comparison of SVGD Bella and VBL. In this section, we compare the performance of SVGD Bella and VBL in distinguishing between correctly and misclassified samples. Specifically, we assess their ability to assign high entropy to misclassified samples and low entropy to correctly classified ones. Figure 11 shows that SVGD Bella outperforms VBL in this regard.
Appendix F Impact of $\gamma$ Values on Robustness.
@@
Appendix G RBF kernel computation in SVGD
The trace term can be written as
$\mathrm{Tr}[\mathbf{A}_{i}\mathbf{A}_{i}^{\top}\mathbf{B}_{i}^{\top}\mathbf{B}_{i}]$ $+\mathrm{Tr}[\mathbf{A}_{i}\mathbf{A}_{i}^{\top}\mathbf{B}_{i}^{\top}\mathbf{B}_{i}]$
$-\mathrm{Tr}[\mathbf{A}_{j}\mathbf{A}_{i}^{\top}\mathbf{B}_{i}^{\top}\mathbf{B}_{j}]-\mathrm{Tr}[\mathbf{A}_{i}\mathbf{A}_{j}^{\top}\mathbf{B}_{j}^{\top}\mathbf{B}_{i}].$
Appendix H Detailed Information on the VQA Task
@@
H.1 Task Definition and Experiment Setup
Task. Visual Question Answering (VQA) (Antol et al. 2015) is a free-form and open-ended task, taking as input an image and a natural-language question about the image and producing a natural-language answer as the output. Questions in VQA require various intelligence capabilities to answer, including image recognition and object detection, as well as reasoning like commonsense reasoning. Below is an overview of the VQA dataset utilized in our paper:
  • VQA v2 (Goyal et al. 2017): This dataset contains 204,721 images, more than 1 Billion (1B) questions, and 10B ground truth answers in total. There are three main types of answers: Yes/No, Number, and Other. The evaluation set contains 80,541, 28,134 and 105,679 questions for Yes/No, Number, and Other respectively.
@@ H.1 Task Definition and Experiment Setup
Models. In our experiments, we applied Bella on top of LLaVA (Liu et al. 2023b, a) to address the VQA task. Specifically, we utilized the LLaVA-1.5 7B model, an improved version of the original LLaVA with superficial modifications. We followed the public training method using a deepspeed codebase.
@@ H.2 Experiments
Models Correct Entropy Incorrect Entropy Accuracy Exact Match
Single base 0.5782 1.4826 70.00 57.17
Ensemble Bella 0.6079 1.5457 67.05 55.26
SVGD Bella 0.5359 1.4602 68.38 56.68
@@
H.3 Results
The performance of different single models and ensembles are presented in Table 4, Table 9. We ensembled two Bella models by simply averaging two softmax output scores. Given the model similarity matrix in Table 10, some observations and conclusions can be made from these tables.
@@ H.3 Results
Figure 12: In comparing LLaVA-VQA predictions on ‘Other’ questions: SVGD Bella outperforms Ensemble Bella, even with more extended predictions. In one case, both Bella err, but with notable human annotator disagreement. Interestingly, SVGD Bella shows greater uncertainty (MI), unlike the confident Ensemble Bella (MI=0). Given that 2 out of 10 annotators chose ”Brown,” SVGD Bella’s answer is deemed acceptable. In another case, SVGD Bella correctly identifies ”Giraffe” albeit singularly, mirroring a minor error also seen among human annotators.
@@
Appendix I Detailed Parameters and Datasets